OpenStudy (anonymous):

A severe weather siren in a local city can be heard within a radius of 1.3 miles. If the mayor of the city wants a new siren that can cover double the area of the old siren, what should the radius of the new siren be? (Round to the nearest tenth of a mile.) a. 3.2 miles b. 2.4 miles c. 2.6 miles d. 1.8 miles

OpenStudy (cornzyblack):

Ok first of all, we find the area that ther 1.5 mile Siren covers using area of a circle, Radius of a circle = 1.3 Area of a circle = pi * r^2 So area = 1.3 * pi = 4.0840 So since the mayor now wants it double that Area it means we multiply the area by 2 So 4.808407 * 2 = 9.6168 However the radius the siren covers has changed, so we are to find the new radius making r subject of formula from area of a circle Area of a circle = 9.618 Area of a circle = pi*r^2 r = \[sqrt{(Area of a circle)/pi} \] r = 1.7495 To the nearest 10th, radius becomes 1.8

OpenStudy (cornzyblack):

well the new radius will give 1.753.... not 1.7495, so if you round up you get 1.8